# UNCOVERing the High-Redshift AGN Population Among Extreme UV Line Emitters

### v1

@@ Abstract
JWST has revealed diverse new populations of high-redshift () AGN and extreme star-forming galaxies that challenge current models. In this paper, we use rest-frame UV emission-line diagnostics to identify AGN candidates and other exceptional ionizing sources, complementing previous studies predominantly focused on broad-line AGN. From a parent sample of 205 UNCOVER galaxies with NIRSpec/PRISM follow-up, we identify 12 C IV, He II, and C III] emitters. Leveraging the combined rest-optical and UV coverage of PRISM, we limit the emission-line model space using the sample’s [O III]/H distribution, significantly decreasing the overlap between AGN and star-formation models in the UV diagnostics. We then find that the five He II emitters are the strongest AGN candidates, with further support from two [Ne V] detections and one X-ray detection from Chandra. We cannot robustly quantify the AGN fraction in this sample, but we note that close to 20% of parent sample galaxies are AGN candidates. The lower-mass line emitters, which are consistent with both AGN and star-forming photoionization models, have more compact sizes and higher specific star formation rates than the parent sample. Higher-resolution and deeper data on these UV line emitters should provide much stronger constraints on the obscured AGN fraction at .
@@ 1 Introduction
In the study of this co-evolution, high-redshift AGN are crucial, placing constraints on SMBH seeding (e.g., Volonteri & Reines, 2016; Dayal et al., 2019; Greene et al., 2020; Inayoshi et al., 2020; Bhowmick et al., 2022; Li et al., 2023; Zhang et al., 2023) as well as AGN feedback and their contribution to reionization (e.g., Fan et al., 2006; Madau & Haardt, 2015; Trebitsch et al., 2023; Dayal et al., 2024; Madau et al., 2024). However, the identification of high-redshift AGN is challenging, partially because low-metallicity stellar populations decrease the efficacy of traditional AGN line ratio diagnostics (e.g., Shapley et al., 2005; Groves et al., 2006; Kewley et al., 2013b, a; Juneau et al., 2014; Steidel et al., 2014; Feltre et al., 2016; Nakajima & Maiolino, 2022; Hirschmann et al., 2019, 2023).
Before the James Webb Space Telescope (JWST, Gardner et al., 2023), the most fruitful approaches to search for high-redshift AGN were UV, X-ray, and radio selection, each with their own biases (for a review, see Fan et al., 2023). From these searches, we knew of a couple hundred AGN at (e.g., Bañados et al., 2016; Wang et al., 2017). However, the handful of pre-JWST AGN all have broad lines (Mortlock et al., 2011; Bañados et al., 2018; Fan et al., 2019; Wang et al., 2021) and spectroscopic follow-up revealed rest-frame UV features comparable to those of the Vanden Berk et al. (2001) lower-redshift broad-line SDSS composite. We are therefore missing a census of the critical population of high-redshift narrow-line AGN. For these objects, which likely comprise the majority of the overall population, there is obscuration of the region of SMBH gravitational influence (i.e., the broad-line region), preventing observations of the AGN continuum and the gas velocity widths (e.g., Hickox & Alexander, 2018).
@@ 3 Line Fitting
In this section, we describe our fitting approach for the UV lines used in the Feltre et al. (2016) diagrams (Section 3.2), H and [O III] (Section 3.3), H and [N II] (Section 3.4), [O II] and [Ne III] (Section 3.3), and the weak, high-ionization line [Ne V] (Section 3.5). We fit the UV lines, H and [O III], and H and [N II] of the entire parent sample but only fit the other lines for our final line emitter sample.
@@ 3.1 Modeling Choices
In each of the fitting regions, we use a power law continuum (normalized with the blue end of the fitting region) with a Gaussian added for each line. We require the lines within each fitting region to share an intrinsic width. We use flat priors on the power law index and amplitude between 3 and 3 and 1e22 and 1e18, respectively. For the fits to relatively weak lines (i.e., not for the +[O III] or +[N II] fits), we simultaneously fit for a multiplicative noise inflation term, using an allowed range of 0.1–10. A factor of 1.4 is typically a good fit. To find best fits and corresponding uncertainties, we use the PyMultiNest (Buchner et al., 2014) implementation of MultiNest (Feroz et al., 2009).
@@
Before calculating likelihoods, we convolve the model with the instrument resolution, using the Jakobsen et al. (2022) JDOX pre-launch instrumental broadening. We adopt a conservative increased resolution of 1.3 (Greene et al., 2024). Point sources have better resolution by a factor closer to 2 (de Graaff et al., 2024; Slob et al., 2024), although this level is not preserved through the reduction and co-adding because the pixel size is large relative to the PSF. We note that lines have a systematic uncertainty on their FWHMs because of our use of this uniform inflation factor rather than a wavelength-dependent, size-dependent factor. To encapsulate this error, for the final emitters, we repeat the fits without any increase on the resolution. We propagate the resulting line flux differences into our plotted ratio and EW errors. This additional source of uncertainty is negligible for all but the highest-SNR sources (45092, 47584, and 45924). Before convolving the instrumental broadening and the model, we use a wavelength grid that oversamples the variable resolution by a factor of four. We then preserve flux while resampling onto the wavelength grid. In Section 5, we present the resulting diagnostic diagrams.
@@
The line fitting presents several challenges: the spectral resolution is both low and variable (R50400), and our understanding of the line spread function (LSF) and noise is incomplete, including regarding the magnitude and impact of correlated noise. Once correlated noise has been quantified for UNCOVER, fitting should use a covariance matrix (as in JADES et al. in prep). To explore the robustness of deblending in light of these uncertainties, we test our fitting procedure on a series of mock spectra. We detail these tests in Appendix A but summarize our main conclusions here. Namely, we demonstrate that we can satisfactorily deblend lines and represent the corresponding flux uncertainties. However, we cannot rule out that a couple of our final sources could be false positives. For one source (34605), of mocks with no injected line yield 3 “detections” and we demonstrate that the false positive rates increase modestly when we use correlated noise in the generation of mocks (e.g., with a correlation coefficient of 0.6 between neighboring pixels).
@@ 3.2 UV Lines
First, we fit C IV 1548, 1551, O III] 1661, 1666, N III] 1750, [Si III] 1883+Si III] 1892, and C III] 1907, 1909, with the fitting region extending to 0.02 rest-frame microns blueward of C IV and redward of C III]. For prior ranges on line fluxes, we use 1e21 up to double the LMFIT (Newville et al., 2016) maximum flux when fitting a power-law continuum plus C IV, He II, and C III]. We require all the lines to share a FWHM between 50 and 1000 km/s. Although the spectral resolution prevents us from constraining the line width, this range allows the fit to marginalize over this source of uncertainty. Because of the blending, we measure the EWs from the model fits.
@@ 3.4 H and [N II]
To check for evidence of broad H, we fit H and the [N II] doublet, with a doublet ratio of 3.05 and the redshift allowed to vary by 0.01 in each direction. Following Greene et al. (2024), for the two-component fit to be preferred over the narrow-only one, we require a 3 improvement on and a 5 detection of the broad component by MultiNest. While there are no new broad-line AGN in our final emitter sample, we report one non-LRD BLAGN, which has evidence for UV lines at the 2.5 level (Figure 14).
@@ 4.1 Models
Individual lines can also serve as AGN indicators when their ionization potential is so high that stars are unlikely to account for any detection. The mere presence of [Ne V] in a galaxy spectrum is often considered sufficient evidence for an AGN, thanks to the ionization potential of 97 eV (e.g., Schmitt, 1998; Gilli et al., 2010; Mignoli et al., 2013), but there are cases of detected [Ne V] from seemingly inactive galaxies (e.g., Izotov et al., 2004, 2012, 2021). Especially at high redshift or low metallicity, [Ne V] may suffer from the same degeneracies as other UV diagnostics. However, the pairing of this line with [Ne III], for instance, helps to distinguish between AGN and other ionization sources (e.g., shocks or Pop III stars, Abel & Satyapal, 2008; Izotov et al., 2021; Cleri et al., 2023; Chisholm et al., 2024). Unfortunately, even in AGN, [Ne V] is intrinsically weak and highly reddened (e.g., Netzer, 1990), so we cannot rely on it for AGN selection, especially at low spectral resolution.
On the other hand, the Feltre et al. (2016) diagnostics are promising, even at the PRISM resolution. To identify these useful line ratios, Feltre et al. (2016) presented AGN models that use CLOUDY (Ferland, 1993; Ferland et al., 1998, 2013) and are parameterized by hydrogen density, ionization parameter, metallicity, dust-to-metal mass ratio of ionized gas, and the power-law index of ionizing spectrum. After pairing the AGN grids with the Gutkin et al. (2016) stellar models, Feltre et al. (2016) found that the key diagnostics leverage the expectation that relatively high values of He II 1640 signal contribution from an AGN narrow-line region (e.g., Villar-Martin et al., 1997; Allen et al., 1998). The success of He II as a tracer is linked to its recombination origin (with an ionization energy of 54.4 eV), whereas the nearby lines are largely collisionally-excited metal lines. We use all of the effective line ratios presented in Feltre et al. (2016) but emphasize the diagrams with C III]/He II vs. C IV/He II and C III]/He II vs. O III]/He II (e.g., Figure 6) because they include the lines that are typically easier to detect.
@@ 4.2 Semi-Empirical Diagnostic Boundaries
We plot the [O III]/H distributions for our parent sample, our line emitters, and the SFG model grids in Figure 4. Using the minimum [O III]/H of 3.0 found among our emitters, we apply a cut to the model grids. In Figure 5 we show the model diagnostic diagram before and after applying this cut. This cut drastically reduces the span of the restricted SFG model grid, such that there is limited overlap between the AGN and SFG models.
For each metallicity and ionization parameter combination, we also plot the fraction of models still included after the cut. The extreme metallicities in the model grids are then disfavored or entirely removed, with the highest fraction of models kept at Z0.005. The effect on ionization parameter, on the other hand, is monotonic: for the stellar models, values of logU lower than cannot survive this cut, and the cut’s effect decreases when moving to higher ionization parameters in the grid. Interestingly, although the full Xiao et al. (2018) model grid extends more than the Gutkin et al. (2016) one into the composite and AGN regions, it occupies less parameter space than the Gutkin et al. (2016) models once the cut has been applied to both. We adopt the updated Gutkin et al. (2016) models in the plotted diagnostics to show the more conservative span of SFG models.
@@ 4.2.2 Observed Line Ratio Diagram Samples from the Literature
First, we note that these diagnostics only apply to the narrow-line region of AGN. The Feltre et al. (2016) AGN models do not consider the temperature and density of the broad-line region and we see that the broad lines of SDSS quasars are deep in the star-forming region. We show this discrepancy using both the SDSS composite (Vanden Berk et al., 2001) and the subset of nitrogen emitters (Batra & Baldwin, 2014). At PRISM resolution in the rest-frame UV, we generally cannot distinguish broad from narrow lines, so we do not expect to select any AGN with UV features dominated by the broad-line region.
Second, the diagnostic works for a variety of narrow-line AGN populations (UV-selected at cosmic noon (Hainline et al., 2011; Alexandroff et al., 2013), optically-selected locally (Nagao et al., 2006), X-ray-selected (Nagao et al., 2006), selected within a sample of C III] emitters (Le Fèvre et al., 2019), and radio galaxies (De Breuck et al., 2000)), but AGN often live in an ambiguous region with model overlap or even just at the edge of the SFG models. Similarly, the star-forming galaxies tend to be consistent with models, but with most being in the Hirschmann et al. (2019) composite region. We include star-forming galaxies and stacks as well as their local low-metallicity dwarf analogues. Despite the benefit of the low-redshift analogues for probing poorly-understood low-metallicity star formation, the few seemingly star-forming objects sitting among the AGN in Figure 6 are all higher-redshift. However, for the star-forming galaxy most consistent with the literature AGN in the log(C III]/He II) vs. log(O III]/He II) diagnostic, the C IV detection is marginal and it is difficult to rule out contribution from an AGN (Amorín et al., 2017). Because of these ambiguities and the PRISM spectral resolution, we focus on identifying the best AGN candidates, noting that these objects would benefit from higher-resolution follow-up.
@@ 5.1 UV Ratio-Ratio Diagnostics
Our final sample of UV line emitters consists of the galaxies with a detection above 5 for C IV or above 3 for He II or C III]. We choose these lines since He II is in every ratio in the diagnostics, C III] is in every y-axis ratio, and C IV (non-)detections are of particular note as we try to separate sources of ionization (Section 4.2.2). Since C IV is unblended, it is easier for a noise feature to be deemed significant at the 3 level; a comparable feature at the location of He II or C III] would split flux with O III] or Si III] in the fit. We therefore use the higher threshold for C IV detections, finding that our suite of mocks also supports this choice (Section A). In Figures 15–17, we show the fixed-redshift and free-redshift fits to this final sample of line emitters. We use the free-redshift fits for the diagnostics to marginalize over uncertainty in the redshift and wavelength solution. From the parent sample of 205 galaxies, we have five He II emitters, eight C III] emitters (four of which are also He II emitters), and seven C IV emitters. Of the C IV emitters, three have only C IV detected.

### v2

@@ Abstract
JWST has revealed diverse new populations of high-redshift () AGN and extreme star-forming galaxies that challenge current photoionization models. In this paper, we use rest-frame UV emission-line diagnostics to identify AGN candidates and other exceptional ionizing sources, complementing previous studies predominantly focused on broad-line AGN. From a parent sample of 205 UNCOVER galaxies with NIRSpec/PRISM follow-up, we identify 12 galaxies with C IV, He II, and/or C III] emission. Three of these galaxies also exhibit clear N III] and/or N IV] lines. Leveraging the combined rest-optical and UV coverage of PRISM, we limit the emission-line model space using the sample’s [O III]/H distribution, significantly decreasing the overlap between AGN and star-formation models in the UV diagnostics. We then find that the five He II emitters are the strongest AGN candidates, with further support from two [Ne V] detections and one X-ray detection from Chandra. Our Balmer line fits also reveal one new broad-line AGN at z=6.87. We cannot robustly quantify the AGN fraction in this sample, but we note that close to 20% of parent sample galaxies are AGN candidates. The lower-mass line emitters, which are consistent with both AGN and star-forming photoionization models, have more compact sizes and higher specific star formation rates than the parent sample. Higher-resolution and deeper data on these UV line emitters should provide much stronger constraints on the obscured AGN fraction at .
@@ 1 Introduction
In the study of this co-evolution, high-redshift AGN are crucial, placing constraints on SMBH seeding (e.g., Volonteri & Reines, 2016; Dayal et al., 2019; Greene et al., 2020; Inayoshi et al., 2020; Bhowmick et al., 2022; Li et al., 2023; Zhang et al., 2023) as well as AGN feedback and their contribution to reionization (e.g., Fan et al., 2006; Madau & Haardt, 2015; Trebitsch et al., 2023; Dayal et al., 2024; Madau et al., 2024). However, the identification of high-redshift AGN is challenging, partially because low-metallicity gas and stellar populations decrease the efficacy of traditional AGN line ratio diagnostics (e.g., Shapley et al., 2005; Groves et al., 2006; Kewley et al., 2013b, a; Juneau et al., 2014; Steidel et al., 2014; Feltre et al., 2016; Nakajima & Maiolino, 2022; Hirschmann et al., 2019, 2023).
Before the James Webb Space Telescope (JWST, Gardner et al., 2023), the most fruitful approaches to search for high-redshift AGN were UV, X-ray, and radio selection, each with their own biases (for a review, see Fan et al., 2023). From these searches, we knew of a couple hundred AGN at (e.g., Bañados et al., 2016; Wang et al., 2017). However, the handful of pre-JWST AGN all have broad lines (Mortlock et al., 2011; Bañados et al., 2018; Fan et al., 2019; Wang et al., 2021) and spectroscopic follow-up revealed rest-frame UV features comparable to those of the Vanden Berk et al. (2001) lower-redshift broad-line SDSS composite. We are therefore missing a census of the critical population of high-redshift narrow-line AGN. For these objects, which likely comprise the majority of the overall population, there is obscuration of the region of SMBH gravitational influence (i.e., the broad-line region), preventing observations of the AGN disk continuum and the gas velocity widths (e.g., Hickox & Alexander, 2018).
@@ 3 Line Fitting
In this section, we describe our fitting approach for the UV lines used in the Feltre et al. (2016) diagrams (Section 3.2), H and [O III] (Section 3.3), H and [N II] (Section 3.4), [O II] and [Ne III] (Section 3.3), and the weak, high-ionization line [Ne V] (Section 3.5). We fit the UV lines, H and [O III], and H and [N II] of the entire parent sample but only fit the other lines for our final sample of galaxies with detected C IV, He II, or C III] emission.
@@ 3.1 Modeling Choices
We separately fit five groups of lines, allowing the continuum to be locally well-fit by a power law. We normalize this power law continuum to the blue end of the fitting region and add a Gaussian for each line. We require the lines within each fitting region to share an intrinsic width. Because we only use line ratios for emission lines that are close in wavelength, we do not model dust. Regardless, especially considering that these UV lines are detectable, reddening effects are small compared to our line ratio uncertainties. Likewise, H absorption should be (e.g., Gil de Paz et al., 2003; Gavazzi et al., 2004), which is negligible compared to the EWs and their uncertainties.
@@
We use flat priors on the power law index and amplitude between 3 and 3 and 1e22 and 1e18, respectively. For the fits to relatively weak lines (i.e., not for the +[O III] or +[N II] fits), we simultaneously fit for a multiplicative noise inflation term, using an allowed range of 0.1–10. A factor of 1.4 is typically a good fit. To find best fits and corresponding uncertainties, we use the PyMultiNest (Buchner et al., 2014) implementation of MultiNest (Feroz et al., 2009).
@@
Before calculating likelihoods, we convolve the model with the instrument resolution, using the Jakobsen et al. (2022) JDOX pre-launch instrumental broadening. We adopt a conservative increased resolution of 1.3 (Greene et al., 2024). Point sources have better resolution by a factor closer to 2 (de Graaff et al., 2024; Slob et al., 2024), although this level is not preserved through the reduction and co-adding because the pixel size is large relative to the PSF. We note that lines have a systematic uncertainty on their FWHMs because of our use of this uniform inflation factor rather than a wavelength-dependent, size-dependent factor. To encapsulate this error, for the final emitters, we repeat the fits without any increase on the resolution. We propagate the resulting line flux differences into our plotted ratio and EW errors. This additional source of uncertainty is negligible (i.e., additional uncertainty on the line ratios) for all but the highest-SNR sources (45092, 47584, and 45924). Before convolving the instrumental broadening and the model, we use a wavelength grid that oversamples the variable resolution by a factor of four. We then preserve flux while resampling onto the wavelength grid. In Section 5, we present the resulting diagnostic diagrams.
@@
The line fitting presents several challenges: the spectral resolution is both low and variable (R50400), and our understanding of the line spread function (LSF) and noise is incomplete, including regarding the magnitude and impact of correlated noise. Once correlated noise has been quantified for UNCOVER, fitting should use a covariance matrix (as in JADES et al. in prep). To explore the robustness of deblending in light of these uncertainties, we test our fitting procedure on a series of mock spectra. We detail these tests in Appendix A but summarize our main conclusions here. Namely, we demonstrate that we can satisfactorily deblend lines and represent the corresponding flux uncertainties. However, we cannot rule out that a couple of our final sources could be false positives. For one source (34605), of mocks with no injected line yield 3 “detections.” We demonstrate that the false positive rates increase modestly when we use correlated noise in the generation of mocks (e.g., with a correlation coefficient of 0.6 between neighboring pixels).
@@ 3.2 UV Lines
First, we fit C IV 1548, 1551, O III] 1661, 1666, N III] 1750, [Si III] 1883+Si III] 1892, and C III] 1907, 1909, with the fitting region extending to 0.02 rest-frame microns blueward of C IV and redward of C III]. Historically, N V 1240 has also been a key AGN indicator (Laporte et al., 2017; Jaskot et al., 2017; Vanzella et al., 2018; Mignoli et al., 2019; Matthee et al., 2022), but it is impossible to isolate with PRISM. For prior ranges on line fluxes, we use from 1e21 to double the LMFIT (Newville et al., 2016) maximum flux when fitting a power-law continuum plus C IV, He II, and C III]. We require all the lines to share a FWHM between 50 and 1000 km/s. Although the spectral resolution prevents us from constraining the line width, this range allows the fit to marginalize over this source of uncertainty. Because of the blending, we measure the EWs from the model fits.
@@ 3.4 H and [N II]
To check for evidence of broad H, we fit H and the [N II] doublet, with a doublet ratio of 3.05 and the redshift allowed to vary by 0.01 in each direction. Following Greene et al. (2024), for the two-component fit to be preferred over the narrow-only one, we require a 3 improvement on and a 5 detection of the broad component by MultiNest. At PRISM resolution, we are only sensitive to broad lines. While there are no new broad-line AGN in our final emitter sample, we report one non-LRD BLAGN, which has evidence for UV lines at the 2.5 level (Figure 14).
@@ 4.1 Models
Individual lines can also serve as AGN indicators when their ionization potential is so high that stars are unlikely to account for any detection. The mere presence of [Ne V] in a galaxy spectrum is often considered sufficient evidence for an AGN, thanks to the ionization potential of 97 eV (e.g., Schmitt, 1998; Gilli et al., 2010; Mignoli et al., 2013), but there are cases of detected [Ne V] from seemingly inactive galaxies (e.g., Izotov et al., 2004, 2012, 2021). Especially at high redshift or low metallicity, [Ne V] may suffer from the same degeneracies as other UV diagnostics. However, the pairing of high-ionization lines with other features (e.g., [Ne III]) helps to distinguish between AGN and other ionization sources (e.g., shocks or Pop III stars, Abel & Satyapal, 2008; Izotov et al., 2021; Cleri et al., 2023; Chisholm et al., 2024; Silcock et al., 2024). Unfortunately, even in AGN, [Ne V] is intrinsically weak and highly reddened (e.g., Netzer, 1990), so we cannot rely on it for AGN selection, especially at low spectral resolution.
On the other hand, the Feltre et al. (2016) diagnostics are promising, even at the PRISM resolution. To identify these useful line ratios, Feltre et al. (2016) presented AGN models that use CLOUDY (Ferland, 1993; Ferland et al., 1998, 2013) and are parameterized by hydrogen density, ionization parameter (i.e., the ratio between the number densities of Hydrogen-ionizing photons and Hydrogen), metallicity, mass ratio of dust and metals, and the power law index of the ionizing spectrum. After pairing the AGN grids with the Gutkin et al. (2016) stellar models, Feltre et al. (2016) found that the key diagnostics leverage the expectation that relatively high values of He II 1640 signal contribution from an AGN narrow-line region (e.g., Villar-Martin et al., 1997; Allen et al., 1998). The success of He II as a tracer is linked to its recombination origin (with an ionization energy of 54.4 eV), whereas the nearby lines are largely collisionally-excited metal lines. We use all of the effective line ratios presented in Feltre et al. (2016) but emphasize the diagrams with C III]/He II vs. C IV/He II and C III]/He II vs. O III]/He II (e.g., Figure 6) because they include the lines that are typically easier to detect.
@@ 4.2 Semi-Empirical Diagnostic Boundaries
We plot the [O III]/H distributions for our parent sample, our line emitters, and the SFG model grids in Figure 4. Using the minimum [O III]/H of 3.0 found among our emitters, we apply a cut to the model grids. In Figure 5 we show the model diagnostic diagram before and after applying this cut. The cut drastically reduces the span of the restricted SFG model grid, such that there is limited overlap between the AGN and SFG models.
For each metallicity and ionization parameter combination, we also plot the fraction of models still included after the cut. Both extremes of metallicity in the model grids are then disfavored or entirely removed, with the highest fraction of models kept at Z0.005. The effect on ionization parameter, on the other hand, is monotonic: for the stellar models, values of logU lower than cannot survive this cut, and the cut’s effect decreases when moving to higher ionization parameters in the grid. Interestingly, although the full Xiao et al. (2018) model grid extends more than the Gutkin et al. (2016) one into the composite and AGN regions, it occupies less parameter space than the Gutkin et al. (2016) models once the cut has been applied to both. We adopt the updated Gutkin et al. (2016) models in the plotted diagnostics to show the more conservative span of SFG models.
@@ 4.2.2 Observed Line Ratio Diagram Samples from the Literature
First, we note that these diagnostics only apply to the narrow-line region of AGN. The Feltre et al. (2016) AGN models do not consider the temperature and density of the broad-line region and we see that the broad lines of SDSS quasars are deep in the star-forming region. We show this discrepancy using both the SDSS composite (Vanden Berk et al., 2001) and the subset of nitrogen emitters (Batra & Baldwin, 2014), as well as with the composite of the subset of Mignoli et al. (2019) C IV-selected AGN that exhibit broad lines. At PRISM resolution in the rest-frame UV, we generally cannot distinguish broad from narrow lines, so we do not expect to select any AGN with UV features dominated by the broad-line region.
Second, the diagnostic works for a variety of narrow-line AGN populations (UV-selected at cosmic noon: Hainline et al. 2011; Alexandroff et al. 2013; Mignoli et al. 2019, optically-selected locally: Nagao et al. 2006, X-ray-selected: Nagao et al. 2006, selected within a sample of C III] emitters: Le Fèvre et al. 2019, and radio galaxies: De Breuck et al. 2000), but AGN often live in an ambiguous region with model overlap or even just at the edge of the SFG models. Similarly, the star-forming galaxies tend to be consistent with models, but with most being in the Hirschmann et al. (2019) composite region. We include star-forming galaxies and stacks as well as their local low-metallicity dwarf analogues. Despite the benefit of the low-redshift analogues for probing poorly-understood low-metallicity star formation, the few seemingly star-forming objects sitting among the AGN in Figure 6 are all higher-redshift. However, for the star-forming galaxy most consistent with the literature AGN in the log(C III]/He II) vs. log(O III]/He II) diagnostic, the C IV detection is marginal and it is difficult to rule out contribution from an AGN (Amorín et al., 2017). Because of these ambiguities and the PRISM spectral resolution, we focus on identifying the best AGN candidates, noting that these objects would benefit from higher-resolution follow-up.
@@ 5.1 UV Ratio-Ratio Diagnostics
Our final sample of UV line emitters consists of the galaxies with a detection above 5 for C IV or above 3 for He II or C III]. We choose these lines since He II is in every ratio in the diagnostics, C III] is in every y-axis ratio, and C IV (non-)detections are of particular note as we try to separate sources of ionization (Section 4.2.2). Since C IV is unblended, it is easier for a noise feature to be deemed significant at the 3 level; a comparable feature at the location of He II or C III] would split flux with O III] or Si III] in the fit. We therefore use the higher threshold for C IV detections, finding that our suite of mocks also supports this choice (Section A). In Figures 15–17, we show the fixed-redshift and free-redshift fits to this final sample of line emitters. We use the free-redshift fits for the diagnostics to marginalize over uncertainty in the redshift and wavelength solution.
From the parent sample of 205 galaxies, we have five He II emitters, eight C III] emitters (four of which are also He II emitters), and seven C IV emitters. Of the C IV emitters, three have only C IV detected.